# FlowSep: Language-Queried Sound Separation with Rectified Flow Matching

### v1

@@ Abstract
Recently, researchers have been exploring separation source systems using non-discriminative models, such as generative models including Generative Adversarial Network (GANs) [17] and diffusion-based models [16]. Due to their generative nature, these models have the potential to enhance the perceptual quality of the separated sources and improve the overall subjective quality. Rectified Flow Matching (RFM) [18] is a recently proposed generative model that models linear relationships between data distributions and noise, providing superior theoretical properties and simplicity. However, it has not yet been explored in sound separation tasks.
@@ II-B1 Rectified Flow Matching
RFM aims to predict the vector field which maps a linear pathway between the noise distribution and target feature . Specifically, RFM learns a neural network to predict the optimal transformation path for any noisy audio feature with a random flow step and parameter . During training, the EFM first computes the noisy version of each data as:
@@ III-A2 Test Set
provides two evaluation sets which can be directly applied for evaluation: DCASE-Synth (DE-S)222https://zenodo.org/records/10886481 and DCASE-Real (DE-R)333https://zenodo.org/records/11425256. In detail, DE-S includes synthetic mixtures mixed from audio clips with an SNR rate ranging from to dB. DE-R consists of audio clips collected from real-world scenarios. Each audio clip contains at least two overlapping sound sources. Each audio clip was annotated their component sources using text descriptions, so that each clip can be used as a mixture from which to extract one or more of the component sources based on a text query. Each audio clip in DE-R was labeled with two such text queries.
@@ III-B Evaluation Metrics
Unlike discriminative networks used in previous systems [13], FlowSep does not separate audio events by masking the mixture input. The separated results are not strictly aligned with the target audio samples in the temporal dimension. Hence, traditional sample-level objective metrics for source separation tasks, such as source-to-distortion ratio (SDR), are not well-suited for evaluating the proposed system. Instead, we assessed its performance using five different metrics. For objective evaluation, we first use frechet audio distance (FAD)[32], a common metric for evaluating generation systems. Next, we apply two metrics based on CLAP [14] including CLAPScore [33], a reference-free metric measuring similarity between the output audio and the text query, and CLAPScoreA, which evaluates the similarity between the output audio and the target audio. For human evaluation, we follow the official methods used in the DCASE2024 Task 9 Challenge, which include assessing the relevance between the target audio and the language query (REL) and the overall sound quality (OVL). Both OVL and REL metrics are rated on a Likert scale from to . All five datasets were evaluated using objective metrics, and we further conducted subjective evaluations on the AudioCaps, DE-S, and DE-R datasets, using a subset of 50 samples randomly selected from each. The subjective evaluations were performed on Amazon Mechanical Turk444https://requester.mturk.com/, with each audio sample rated by at least different raters.
@@ IV-B Experimental Details
We first train a kHz BigVGAN [23] on the training datasets. For the encoder and decoder of the FlowSep, we apply the FLAN-T5-large [19] checkpoint and the pre-trained VAE model from AudioLDM [27] model. The RFM model is trained for M steps with a batch size of and a base learning rate of . We use the publicly released LASS-Net [7] and AudioSep [13] models as baselines. For the diffusion-based network, denoted as DiffusionSep, we replace the RFM with diffusion loss [27] and train the system using the same parameters as mentioned above.
@@ IV-D Case Studies
We visualized the separated spectrograms of AudioSep and FlowSep using an example from the DE-S testing set. As shown in Figure 3, The spectrogram generated by FlowSep is closely matching the ground truth. In contrast, AudioSep’s results show incomplete separation with noticeable spectral gaps. The results highlight that, unlike previous discriminative approaches, FlowSep, as a generative network, demonstrates promising capabilities in source separation tasks. More case studies can be found online555https://audio-agi.github.io/FlowSep_demo.
@@ IV-E Efficiency Analysis
We conduct an efficiency analysis of FlowSep in comparison with AudioSep and a diffusion-based LASS model we implemented, with the performance results presented in Table III. The inference experiments were conducted on the AudioCaps dataset using a single A100 GPU with 80 GB of memory. As compared with AudioSep, our generative approach does not present better efficiency than discriminative approaches, while FlowSep generates results with better quality. Furthermore, FlowSep surpasses the diffusion-based LASS model in both FAD and CLAPScore metrics with the same number of inference steps (e.g., steps). FlowSep-RFM maintains stable performance when the inference steps are reduced (e.g., only minor degradation when reducing the steps from to ) and achieves acceptable performance with even fewer steps (e.g., 10 steps). These findings underscore that RFM outperforms diffusion-based methods in separation tasks, effectively reducing inference time while enhancing separation performance.

### v2

@@ Abstract
Recently, researchers have been exploring separation source systems using non-discriminative models, such as generative models including Generative Adversarial Network (GANs) [17] and diffusion-based models [16]. Due to their generative nature, these models have the potential to enhance the perceptual quality of the separated sources and improve the overall subjective quality. Rectified Flow Matching (RFM) [18] is a recently proposed generative model. Similar to diffusion-based systems that learn to gradually denoise the output from general distribution, RFM models linear the relationships between data distributions and noise and provide superior theoretical properties and simplicity. However, it has not yet been explored in sound separation tasks.
@@ II-B1 Rectified Flow Matching
RFM aims to predict the vector field which maps a linear pathway between the noise distribution and target feature . Specifically, RFM learns this UNet neural network to predict the optimal transformation path for any noisy audio feature with a random flow step and parameter . During training, the RFM first computes the noisy version of each data as:
@@ III-A2 Test Set
provides two evaluation sets which can be directly applied for evaluation: DCASE-Synth (DE-S) and DCASE-Real (DE-R). In detail, DE-S includes synthetic mixtures mixed from audio clips with an SNR rate ranging from to dB. DE-R consists of audio clips collected from real-world scenarios. Each audio clip contains at least two overlapping sound sources. Each audio clip was annotated their component sources using text descriptions, so that each clip can be used as a mixture from which to extract one or more of the component sources based on a text query. Each audio clip in DE-R was labeled with two such text queries.
@@ III-B Evaluation Metrics
Unlike discriminative networks used in previous systems [13], FlowSep does not separate audio events by masking the mixture input. The separated results are not strictly aligned with the target audio samples in the temporal dimension. Hence, traditional sample-level objective metrics for source separation tasks, such as source-to-distortion ratio (SDR), are not well-suited for evaluating the proposed system. Instead, we assessed its performance using five different metrics. For objective evaluation, we first use frechet audio distance (FAD)[32], a common metric for evaluating generation systems. Next, we apply two metrics based on CLAP [14] including CLAPScore [33], a reference-free metric measuring similarity between the output audio and the text query, and CLAPScoreA, which evaluates the similarity between the output audio and the target audio. For human evaluation, we follow the official methods used in the DCASE2024 Task 9 Challenge, which include assessing the relevance between the target audio and the language query (REL) and the overall sound quality (OVL). Both OVL and REL metrics are rated on a Likert scale from to . All five datasets were evaluated using objective metrics, and we further conducted subjective evaluations on the AudioCaps, DE-S, and DE-R datasets, using a subset of 50 samples randomly selected from each. The subjective evaluations were performed on ten different listeners, with six researchers in audio and speech and four from other fields.
@@ IV-B Experimental Details
We first train a kHz BigVGAN [23] on the training datasets. For the encoder and decoder of the FlowSep, we apply the FLAN-T5-large [19] checkpoint and the pre-trained VAE model from AudioLDM [27] model. We freeze all the other components and the RFM model is trained for M steps with a batch size of and a base learning rate of . We use the publicly released LASS-Net [7] and AudioSep [13] models as baselines. For the diffusion-based network, denoted as DiffusionSep, we replace the RFM with general diffusion-based loss [27] and train the system using the same parameters.
@@ IV-D Case Studies
We visualized the separated spectrograms of AudioSep and FlowSep using an example from the DE-S testing set. As shown in Figure 3, the spectrogram generated by FlowSep is closely matching the ground truth. In contrast, AudioSep’s results show incomplete separation with noticeable spectral gaps. The results highlight that, unlike previous discriminative approaches, FlowSep, as a generative network, demonstrates promising capabilities in source separation tasks. More case studies can be found online222https://audio-agi.github.io/FlowSep_demo.
@@ IV-E Efficiency Analysis
We conduct an efficiency analysis of FlowSep in comparison with AudioSep and a diffusion-based LASS model we implemented, with the performance results presented in Table III. The inference experiments were conducted on the AudioCaps dataset using a single A100 GPU with 80 GB of memory. As compared with AudioSep, our generative approach does not present better efficiency than discriminative approaches, while FlowSep generates results with better quality. Furthermore, FlowSep surpasses the diffusion-based LASS model in both FAD and CLAPScore metrics with the same number of inference steps (e.g., steps). FlowSep-RFM maintains stable performance when the inference steps are reduced (e.g., only minor degradation when reducing the steps from to ) and achieves acceptable performance with even fewer steps (e.g., 10 steps). These findings underscore that RFM outperforms diffusion-based methods in separation tasks, and our proposed system can effectively reducing inference time while enhancing separation performance.